Bismuth-213 is used in radiation therapy to treat leukemia, while strontium-89 is utilized to treat bone metastasis. Write the isotopic symbols of the two atoms. How many protons and neutrons does each atom have?

1
Expert's answer
2022-02-10T04:55:05-0500

Solution:

A = mass number

Z = atomic number (number of protons)

N = number of neutrons


A = Z + N


(1):

Bismuth-213 = 213Bi

So 213Bi has a mass number (A) of 213, which is the sum of the protons (Z) and neutrons (N) in the nuclei of that isotope's atoms.

The number of protons (Z) is the atomic number of bismuth (Z = 83; according to the periodic table).

So the number of neutrons is: N = A - Z = 213 - 83 = 130.

Answer (1): Bismuth-213: Z = 83; N = 130.


(2):

Strontium-89 = 89Sr

So 89Sr has a mass number (A) of 89, which is the sum of the protons (Z) and neutrons (N) in the nuclei of that isotope's atoms.

The number of protons (Z) is the atomic number of strontium (Z = 38; according to the periodic table).

So the number of neutrons is: N = A - Z = 89 - 38 = 51.

Answer (2): Strontium-89: Z = 38; N = 51.